### v2

# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

Abstract
Recent advancements in machine learning, particularly in Natural Language Processing (NLP), have led to the development of sophisticated models trained on extensive datasets, yet raising concerns about the potential leakage of sensitive information. In response, regulatory measures such as the European Union’s General Data Protection Regulation (GDPR) have driven increasing interest in Machine Unlearning techniques, which enable models to selectively forget specific data entries. Early approaches primarily relied on pre-processing methods, while more recent research has shifted towards training-based unlearning techniques. Despite their effectiveness, most existing methods require access to the original training data, which is often inaccessible. Additionally, directly applying unlearning techniques bear the cost of undermining the model’s expressive capabilities. To address these challenges, we introduce the Iterative Contrastive Unlearning (ICU) framework, which consists of three core components: A Knowledge Unlearning Induction module designed to remove specific knowledge through an unlearning loss; A Contrastive Learning Enhancement module to preserve the model’s expressive capabilities against the pure unlearning goal; And an Iterative Unlearning Refinement module that dynamically assess the unlearning extent on specific data pieces and make iterative update. Experimental results demonstrate the efficacy of our ICU method in unlearning sensitive information while maintaining the model’s overall performance, offering a promising solution for privacy-conscious machine learning applications.
1 Introduction
With the continuous evolution of machine learning, we have witnessed an unprecedented expansion in the size of models and the diversity of data used for their training. Particularly, the field of Natural Language Processing (NLP) has experienced remarkable progress, driven by the development of advanced Generative Language Models (GLMs) such as GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and Google Gemini (Team et al., 2023). Despite these advancements, concurrent studies have highlighted a critical concern: the potential leakage of sensitive information (Carlini et al., 2021), including phone numbers, email addresses, and other personal data embedded within the training datasets (Brown et al., 2022; Carlini et al., 2022). In response to these privacy concerns, regulations such as the General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) have codified the “Right To Be Forgotten” (RTBF) (Mantelero, 2013; Villaronga et al., 2018), leading to the exploration of Machine Unlearning techniques. These techniques aim to make models effectively “forget” specific data, treating them as if they were never part of the training dataset.
In the initial stages of research, various pre-processing methods were proposed to achieve unlearning on certain data. For instance, Kandpal et al. (2022) found that the likelihood of language models regenerating training sequences is correlated with the frequency of those sequences in the training set. They showed that deduplicating the training data makes language models significantly more resilient against privacy attacks. However, such methods are often time-consuming and resource-intensive, making them impractical for scenarios with frequent unlearning requests.
More recently, researchers have shifted their focus to training-based machine unlearning approaches, which modify the training process itself rather than solely manipulating the data. For example, SISA (Bourtoule et al., 2021) partitions the original dataset into several non-overlapping shards and then aggregates models trained on these separate shards. When handling data deletion requests, only the models trained on the affected shards need to be retrained. KGA (Wang et al., 2023) introduces an additional dataset, using it to fine-tune the original model alongside the original dataset. Nonetheless, both of these methods assume that the training data remains accessible during the unlearning process. In practice, the training data for large language models may not be available after model deployment, rendering such methods infeasible. A recent method, Knowledge Unlearning for Mitigating Privacy Risks (KUMPR) (Jang et al., 2023), was designed to address this scenario.111We refer to this method as KUMPR as the authors did not provide a specific name. KUMPR reverses the training objective by maximizing the negative log-likelihood for target tokens and uses metrics to determine whether the model has “forgotten” a target sequence.
However, directly reversing the training objective of language models can undermine their overall expressive capabilities, beyond merely forgetting the intended information. As illustrated in Figure 1, given a training sample with the prefix “Who is Harry Potter?” and its reference suffix, the goal is to make the model “forget” specific knowledge, such as “J.K. Rowling” and “1997”. Although the existing methods (e.g., KUMPR) successfully forget the key information contained in the reference suffix, the resulting model loses its general expressive ability, often generating repetitive, nonsensical text like “I have a Harry Potter”. To address this limitation, we propose an Iterative Contrastive Unlearning (ICU) framework, which aims to achieve unlearning while preserving the model’s overall generalization ability.
More specifically, our ICU framework comprises three components: (1) a Knowledge Unlearning Induction (KUI) module, which targets specific knowledge to be unlearned from the model; (2) a Contrastive Learning Enhancement (CLE) module, which samples paired data from the analogous documents of target knowledge. Subsequently, we further design two learning enhancement losses to maintain the generalization ability against the unlearning process. (3) an Iterative Unlearning Refinement (IUR) module, which assess the unlearning extent on specific data pieces and updates the unlearning dataset dynamically. In brief, our contributions are as follows:
We conduct an in-depth study on the unlearning techniques for Generative Language Models, specifically focusing on maintaining the expression ability while effectively unlearning, an area that has been largely overlooked in previous research.
We propose the ICU framework, which consists of three components: Knowledge Unlearning Induction, Contrastive Learning Enhancement, and Iterative Unlearning Refinement.
We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/himalalps/ICU.
2 Related work
2.1 Machine unlearning
Machine unlearning, introduced by Cao & Yang (2015), aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data was never part of the training set. Traditional approaches (Nguyen et al., 2022) that exclude specific data from training datasets and retrain the model are highly time-consuming and resource-intensive, making them impractical for modern deep neural networks. Similar methods involving retraining (Bourtoule et al., 2021; Kumar et al., 2023) also struggle with scalability, particularly when handling numerous deletion requests or when comprehensive datasets are not readily available. To address these challenges, researchers have explored approximate unlearning techniques (Golatkar et al., 2020; Guo et al., 2019; Mehta et al., 2022). One such approach involves data pre-processing, which efficiently identifies and removes sensitive information before model training. Kandpal et al. (2022) applied this method to structured private data, such as phone numbers and medical records, and found it effective. However, challenges arise when dealing with unstructured data, as pre-processing may not fully remove all sensitive information (Brown et al., 2020) and cannot comprehensively address ongoing deletion demands (Brown et al., 2022).
Recent studies (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023) have focused on fine-tuning language models to tackle machine unlearning challenges. Jang et al. (2023) proposed a novel approach by reversing the traditional training objective, aiming to maximize rather than minimize the negative log-likelihood of tokens designated for forgetting. Despite effectiveness, this kind of methods cannot avoid undermining models’ generalization ability. Other recent methods (Chen & Yang, 2023; Gao et al., 2024; Gu et al., 2024; Kassem et al., 2023; Wang et al., 2023; Maini et al., 2024; Yao et al., 2023) use diverse techniques such as knowledge gap alignment and reinforcement learning. Despite unlearning effectively, these methods are often complex and computationally expensive, limiting their practicality. For instance, Gu et al. (2024) utilized second-order information (Hessian) to provide stronger guarantees for data removal while maintaining model utility, but this approach requires substantial computational resources for Hessian approximation, making it difficult to play a role in real scenarios.
2.2 Language models
Language models are designed to understand, generate, and predict human language (Zhao et al., 2023), which have gained considerable attention in recent years (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023).
Traditional language models, such as rule-based approaches (Weizenbaum, 1966) and statistical models (Brown et al., 1992), generate outputs that resemble human language but do not perfectly reflect the training data. Early neural network models in NLP, including Recurrent Neural Networks (RNNs) (Werbos, 1990), faced limitations due to their sequential processing architecture, which resulted in high computational demands and hindered scalability.
The introduction of the transformer architecture (Vaswani et al., 2017) revolutionized NLP by enabling the effective capture of contextual relationships through self-attention mechanisms. Subsequent advancements have led to the development of three primary categories of transformer-based models: encoder-only models, such as BERT (Devlin et al., 2018); decoder-only models, exemplified by GPT (Radford et al., 2018); and encoder-decoder models, like T5 (Raffel et al., 2020). Decoder-only models, including GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and others (Anil et al., 2023; Team et al., 2023; Touvron et al., 2023), have demonstrated exceptional performance across a wide range of NLP tasks. However, this success has raised privacy concerns due to the potential leakage of sensitive information from the training data. Additionally, increasing model capacity has led to greater demands for training data and computational resources (Achiam et al., 2023; Anthropic, 2024; Hoffmann et al., 2022; Ouyang et al., 2022), creating significant challenges for researchers working on machine unlearning for advanced models.
3 ICU framework
3.1 Problem statement
Given the data and model , where is the collection of pieces of data to be forgotten and is the original language model with its parameters denoted as , machine unlearning aims to modify the parameters such that the retention of previously learned information about is minimized while maintaining desirable model performance and meeting specified constraints.
3.2 Model overview
We propose a novel Iterative Contrastive Unlearning (ICU) framework, illustrated in Figure 2. This framework focuses on unlearning for decoder-only models, addressing the challenge of mitigating the memorization of sensitive information while preserving language generation capabilities. In addition to (a) Knowledge Unlearning Induction module, which trains the model to forget target sequences, we introduce two supplementary modules. (b) Contrastive Learning Enhancement module utilizes specially selected data to maintain overall model performance during unlearning. Furthermore, (c) Iterative Unlearning Refinement module updates the data to be forgotten in an iterative manner, preventing over-unlearning and mitigating performance degradation.
3.3 Knowledge Unlearning Induction
For a sample in the forget set , the sequence of tokens is denoted as . Following the approach of Jang et al. (2023), we negate the original negative log-likelihood of the target token sequences to induce the model to forget these sequences. The unlearning objective is defined as:
where denotes the first tokens of the sequence, and represents the conditional probability of predicting the next token given the previous tokens , with representing the model parameters.
3.4 Contrastive Learning Enhancement
To maintain the model’s stable generation capabilities during unlearning, we propose training the model simultaneously on analogous data. This approach ensures that the model forgets specific information without significantly reducing its ability to recognize and generate similar patterns.
Analogous data Construction.
The first step is to construct a data pool related to the forget set , aiming to identify data samples that contain similar but different knowledge. Specifically, we retrieve documents from Wiki that belong to the same category as the forget set but contain different key concepts. This process results in the creation of the Analogous Set , which will be used in subsequent steps.
KNN Sampling.
In this part, we compute the sentence embeddings for all samples in using a pre-trained sentence transformer . For each sample in with its embedding , we employ K-Nearest Neighbors (KNN) (Douze et al., 2024) to identify the nearest () embedding and the corresponding sample . All are then collected to form :
where is the cosine similarity function used in KNN search.
Learning Enhancement.
The retrieved paired data for is denoted as . In contrast to the unlearning objective, we force the model to learn patterns from these paired token sequences using the negative log-likelihood, defined as follows:
3.5 Iterative Unlearning Refinement
Unlike conventional machine learning techniques, validating the efficacy of unlearning presents challenges in identifying a suitable validation set, as is integrated into the training phase. Thus, establishing an appropriate stopping criterion is essential. After each training epoch, the model’s performance relative to the target data is evaluated using the metrics described below:
where and are predefined thresholds for the iteration process. We empirically determine that a specific token sequence is considered “forgotten” if Equation (5) is satisfied. Samples deemed “forgotten” are excluded from subsequent epochs. This iterative refinement process serves a dual purpose: signaling the end of training and preventing the unnecessary erosion of already discarded information, thus preserving the model’s proficiency and effectiveness.
Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) is a metric originally used to evaluate machine translation quality by measuring the similarity between a model-generated token sequence and one or more reference translations, based on n-gram comparisons. The BLEU score ranges from 0 to 1, with a higher score indicating a better match to the reference.
BERTScore (Zhang et al., 2019) leverages contextual embeddings from BERT (Devlin et al., 2018) models to assess the similarity between two provided sentences. Unlike previous metrics that rely solely on exact word n-grams, BERTScore considers semantic similarity, offering a more adaptable and accurate measure of sentence similarity.
3.6 Training
The objectives in Equation (1), (3) and (4) are jointly used to optimize the model during unlearning. The training process is governed by minimizing the following loss function:
where are positive hyper-parameters that control the contributions of the different optimizing objectives.
4 Experiments
In this section, we first describe the datasets used for training and evaluation, as well as the metrics employed to assess performance. Next, we introduce the baseline methods used for comparison with our proposed approach, followed by the configuration details of our method. Finally, we present and analyze the experimental results.
4.1 Datasets
To evaluate ICU’s learning and unlearning capabilities, we selected two types of datasets: Target Datasets, which assess the unlearning performance, and Downstream Dataset, which evaluates the original capabilities of the models.
Target Dataset.
The Pile corpus (825GB) is a large dataset constructed from 22 diverse high-quality subsets, many of which derive from academic or professional sources (e.g. books, open source code) (Gao et al., 2020). As the whole dataset is not available at present, we use a subset of the Pile corpus, which is released as a benchmark for data extraction attacks.222https://github.com/google-research/lm-extraction-benchmark Designed to be easy-to-extract, the subset contains 15,000 samples, randomly sampled from the Pile training dataset. Most of them are in English, but there are also samples in Russian or Chinese. Each sample consists of a 200-token sequence, among which are 100 pre-prefix tokens, 50 prefix tokens, and 50 suffix tokens. Following Jang et al. (2023), we only use the prefix and suffix tokens in our experiments.
Downstream Dataset.
To assess the general performance of the LMs subsequent to the process of unlearning, a diverse array of downstream tasks is employed. This endeavor is aimed at ensuring that the original capabilities of the models remain unaffected.
This evaluation encompasses nine distinct classification tasks spanning three thematic domains. Specifically, these domains include linguistic reasoning tasks such as Hellaswag (Zellers et al., 2019) and Lambada (Paperno et al., 2016), as well as assessments of commonsense reasoning through Winogrande (Sakaguchi et al., 2021) and COPA (Gordon et al., 2012). Additionally, scientific reasoning abilities are evaluated through tasks such as ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), Piqa (Bisk et al., 2020), MathQA (Amini et al., 2019), and PubmedQA (Jin et al., 2019).
Furthermore, four dialogue tasks, namely Wizard of Wikipedia (Dinan et al., 2018), Empathetic Dialogues (Rashkin et al., 2018), Blended Skill Talk (Smith et al., 2020), and Wizard of Internet (Komeili et al., 2021), are used to gauge the model’s proficiency in generating coherent responses. In addition, we measure the perplexity of the unlearned models on the validation set of Pile and Wikitext. Following (Jang et al., 2023), we use the test set for Lambada and the validation set for the remaining tasks.
4.2 Metrics
As stated in Section 4.1, we assess both learning and unlearning capabilities using two types of datasets. For evaluating unlearning performance, we follow Jang et al. (2023) and examine the forgetting effect on the target unlearning data using Extraction Likelihood (EL) and Memorization Accuracy (MA). In the work of Jang et al. (2023), these metrics are also used as stopping criteria during training, making them unsuitable as sole evaluation metrics. As mentioned in Section 3.5, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.
To evaluate the original capabilities of unlearned models, we first use the basic metrics provided in the downstream datasets, obtaining Accuracy for classification task and F1 for dialogue tasks. Besides, We adopt Information Entropy to measure the expression performance in the results. Furthermore, we also employ GPT-4 and Human Evaluation to assess the text generated by the unlearned models. The following details the metrics used:
Extraction Likelihood (EL) is introduced by Jang et al. (2023) to measure the average success rate of varying extraction attacks quantified via getting the n-gram overlap of generated and target token sequences. It is computed by the following equation:
Memorization Accuracy (MA) (Tirumala et al., 2022) quantifies how much model has memorized the given token sequences, which is defined as follows:
Information Entropy quantifies the average uncertainty in a set of outcomes, reflecting the amount of information produced by a random source. Higher entropy indicates greater unpredictability and information content. Mathematically, entropy () is defined for a discrete random variable with possible outcomes and corresponding probabilities as:
GPT Evaluation uses GPT-4 (Achiam et al., 2023) to evaluate the unlearned models in two perspectives: whether the model generates text without prior knowledge of key information in the referenced target data and whether the generated sequences are coherent. The prompts can be found in Appendix A.
Human Evaluation indicates that we hire human experts to determine the goodness of the generated response, following the goal in the GPT Evaluation part. The annotation agreement can be found in Appendix B.
4.3 Baseline methods
Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) (Black et al., 2021), which is pre-trained on the Pile corpus. Following Jang et al. (2023), we utilize the Opt model family (125M, 1.3B, 2.7B) (Zhang et al., 2022), which is pre-trained on a deduplicated version of the Pile as well as other corpus, serving as our baseline method for deduplication since the deduplicated version of GPT-Neo by Kandpal et al. (2022) is not publicly available. For the approximate unlearning methods, we include KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024) and LLMU (Yao et al., 2023) as other baseline methods on GPT-Neo models to show the effectiveness of our proposed method. We follow their publicly released codes and the same training and evaluation procedure to obtain the results.
4.4 Configurations
For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples. In the Contrastive Learning Enhancement module (Section 3.4), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling. The model is optimized by Adam (Kingma & Ba, 2014) with a learning rate of , and . We regard the model to have “forgotten” the target dataset with an average of and . The filtering thresholds during iteration are and as introduced in Section 3.5.
We run all the experiments on a Linux server with one 2.60GHz Intel Xeon Platinum 8358 CPU and NVIDIA GeForce RTX 3090 GPUs. We use one GPU for 125M models with batch size of 8. With Deepspeed Stage 2, we use three for 1.3B and six for 2.7B respectively with batch size of 4.
Model | # | EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | GPT | Epoch |
Params | (%) | (%) | (F1) | (ACC) | (F1) | (PPL) | (PPL) | ||||
Neo (before unlearning) | 125M | 51.9 | 76.8 | 70.3 | 4.139 | 43.5 | 10.0 | 20.1 | 38.0 | - | - |
Opt | 7.5 | 52.9 | 49.2 | 3.014 | 42.7 | 10.8 | 29.1 | 38.0 | 3.08 | - | |
Neo + KUMPR | 0.7 | 19.1 | 29.7 | 0.712 | 35.1 | 3.7 | >10000 | >10000 | 1.05 | 3.2 | |
Neo + DPO | 17.4 | 49.4 | 42.2 | 1.643 | 39.0 | 1.8 | 61.2 | 158.5 | 1.99 | 79.0 | |
Neo + KL | 4.9 | 56.2 | 54.3 | 3.670 | 42.6 | 9.9 | 27.0 | 54.0 | 3.64 | 14.0 | |
Neo + LLMU | 3.3 | 58.7 | 43.1 | 1.825 | 42.7 | 10.3 | 23.6 | 47.4 | 2.98 | 7.0 | |
Neo + ICU (ours) | 4.4 | 55.6 | 53.3 | 3.833 | 43.3 | 10.3 | 21.6 | 40.1 | 3.92 | 21.4 | |
Neo (before unlearning) | 1.3B | 98.2 | 92.3 | 86.3 | 4.640 | 49.7 | 12.3 | 13.2 | 18.7 | - | - |
Opt | 31.0 | 67.8 | 65.8 | 3.856 | 51.7 | 13.3 | 18.0 | 19.2 | 3.63 | - | |
Neo + KUMPR | 0.8 | 8.1 | 26.6 | 0.817 | 34.0 | 0.1 | >10000 | >10000 | 1.26 | 2.0 | |
Neo + DPO | 20.0 | 58.4 | 55.1 | 2.564 | 44.8 | 4.8 | 26.6 | 44.6 | 2.56 | 30.0 | |
Neo + KL | 3.7 | 61.5 | 38.7 | 1.627 | 47.4 | 11.8 | 17.5 | 26.0 | 2.18 | 7.0 | |
Neo + LLMU | 4.4 | 63.3 | 41.4 | 1.805 | 47.4 | 11.9 | 17.5 | 25.7 | 2.43 | 8.0 | |
Neo + ICU (ours) | 4.7 | 51.3 | 52.7 | 3.900 | 49.0 | 12.1 | 14.1 | 19.3 | 4.33 | 29.2 | |
Neo (before unlearning) | 2.7B | 96.7 | 93.7 | 90.2 | 4.719 | 52.4 | 12.3 | 12.0 | 16.2 | - | - |
Opt | 34.4 | 70.1 | 66.8 | 3.921 | 53.9 | 13.7 | 16.3 | 16.7 | 3.65 | - | |
Neo + KUMPR | 1.4 | 18.7 | 26.5 | 0.519 | 34.0 | 5.4 | 4926.2 | >10000 | 1.00 | 6.6 | |
Neo + DPO | 20.4 | 58.6 | 57.5 | 2.772 | 49.0 | 7.6 | 24.1 | 36.5 | 2.66 | 31.0 | |
Neo + KL | 3.4 | 62.5 | 41.3 | 1.730 | 51.3 | 12.5 | 14.7 | 20.9 | 2.39 | 6.0 | |
Neo + LLMU | 4.9 | 63.2 | 40.7 | 1.639 | 50.9 | 12.4 | 16.0 | 22.4 | 2.26 | 8.0 | |
Neo + ICU (ours) | 4.5 | 48.3 | 52.7 | 3.725 | 52.1 | 12.1 | 13.1 | 17.0 | 4.40 | 32.6 |
4.5 Main results
The results of all methods are summarized in Table 1. Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines. Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the generation ability of the model. Compared to Opt, our model unlearns the original model better. In terms of GPT Evaluation, our model achieves the best score compared to all baselines. Additionally, we observe several interesting phenomena:
First, the impact of unlearning becomes more pronounced with larger models, indicating that larger models have a higher tendency to memorize sensitive information, which our method effectively mitigates. Second, while KUMPR significantly forgets sensitive information, it also impairs the model’s general performance (e.g., PPL). In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method. Full results for each sample are provided in Appendix G.
4.6 Ablation study
As discussed in Section 3.4, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability. To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively. The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity. Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
4.7 Parameter sensitivity
EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | Epoch | ||
(%) | (%) | (F1) | (ACC) | (F1) | (PPL) | (PPL) | ||||
0.1 | 0.5 | 3.6 | 54.7 | 47.0 | 3.009 | 39.5 | 5.6 | 95.2 | 212.7 | 9.4 |
0.5 | 0.1 | 3.3 | 56.5 | 49.9 | 3.423 | 40.0 | 8.7 | 35.4 | 85.5 | 9.4 |
0.5 | 0.5 | 4.2 | 57.5 | 52.6 | 3.755 | 42.4 | 9.7 | 24.5 | 48.8 | 14.4 |
0.5 | 1.0 | 4.4 | 55.6 | 53.3 | 3.833 | 43.3 | 10.3 | 21.6 | 40.1 | 21.4 |
1.0 | 0.5 | 4.5 | 55.1 | 52.1 | 3.700 | 43.1 | 10.1 | 21.6 | 40.1 | 22.6 |
1.0 | 1.0 | 4.6 | 53.5 | 52.7 | 3.846 | 43.2 | 10.0 | 21.6 | 40.0 | 27.2 |
1.0 | 2.0 | 4.5 | 47.7 | 50.6 | 3.686 | 43.1 | 10.0 | 21.6 | 39.7 | 37.4 |
2.0 | 1.0 | 4.6 | 47.2 | 50.6 | 3.809 | 42.7 | 9.8 | 22.1 | 40.5 | 40.8 |
2.0 | 2.0 | 4.5 | 45.8 | 51.2 | 3.843 | 42.9 | 10.0 | 22.1 | 40.1 | 45.4 |
To examine the influence of the loss hyperparameters and in Section 3.4, we conducted extensive parameter sensitivity experiments on GPT-Neo 125M. The results are summarized in Table 2.
In general, increasing and enhances learning ability while diminishing unlearning ability, with the exception of memorization accuracy (MA). For MA, which assesses the model’s memory capacity, the model effectively memorizes corresponding tokens through paired data, thereby retaining its original generative capability. Meanwhile, we find that when the learning weight is less than the unlearning weight of 1, variations in the regularization weight significantly impact the model’s performance. Conversely, when exceeds the forgetting weight of 1, changes in do not significantly affect performance. This indicates that both hyperparameters contribute to increased learning ability, corroborating the findings presented in Figure 3. To balance the model’s learning and forgetting abilities, we ultimately selected and as our reported parameters.
4.8 Case study
To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning. As shown in Figure 4, the reference includes sensitive information such as “@ThingsExpo 2016”. Before unlearning, the original models (e.g., GPT-Neo and Opt) retain and reproduce this information when prompted with the corresponding prefix. When applying the KUMPR method, the models lose their original conversational abilities and repetitively output the word “the”. DPO method also produces sentences not coherent. KL and LLMU methods both repeat same sentence. In contrast, our approach effectively forgets the sensitive information while learning the correct outputs from paired data, preserving the model’s generative capabilities. This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework. Additional cases are provided in Appendix F.
More experimental analyses, such as Human Evaluation and Generalization Ability of ICU, can be found in Appendix A - G.
5 Conclusion and future work
In this paper, we explored the direction of machine unlearning for Generative Language Models. We analyzed the challenges inherent in this topic and proposed an Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. Extensive experiments on models of three different sizes demonstrated the effectiveness of our proposed method.
For future work, we plan to explore the application of our framework to multi-modal data, extending unlearning capabilities to models that integrate text, images, and other data types. Additionally, we are very interested in studying the implications of unlearning algorithms in privacy-sensitive real-world applications, such as healthcare and personalized services.
Appendix A GPT prompts
The prompts we employed in the evaluation part (Section 4.2) is as follows:
Given a prefix and a reference suffix, candidates will generate text based on the prefix. Your task is to evaluate the differences between each candidate’s text and the provided reference. The candidate should act as if he has no prior knowledge of the reference. The final score out of 10 should be in final line in the format of "Finale score: X".
Example:
Prefix:
"Who is Harry Potter?"
Reference:
"Harry Potter is a fictional character created by British author J. K. Rowling."
Candidate:
"Harry Potter is a character in American comics."
Evaluation: The candidate’s response is different from the reference but makes sense if the candidate does not know Harry Potter. The response is fluent and coherent.
Final score: 10
Example:
Prefix:
"Who is Harry Potter?"
Reference:
"Harry Potter is a fictional character created by British author J. K. Rowling."
Candidate:
"Harry Potter is a character in a novel series written by J. K. Rowling."
Evaluation: The candidate’s response is similar to the reference so should be rated lower.
Final score: 0
Example:
Prefix:
"Who is Harry Potter?"
Reference:
"Harry Potter is a fictional character created by British author J. K. Rowling."
Candidate:
"Harry Potter is Harry Potter is Harry Potter is
Evaluation: The candidate’s response is repetitive and lacks coherence.
Final score: 2
Now it’s your turn:
Prefix: "<Prefix>"
Reference: "<Reference>"
Candidate: "<Candidate>"
Appendix B Human evaluation
To enhance the credibility of the GPT-4 grading, we also include human annotation results. Specifically, for each sample, human evaluators rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference). The annotators are only presented with only the prefix, referenced suffix and the generated text by the model and are unaware of the full scope of the work.
The comparison between GPT and Human Evaluation scores can be found in Table 3. The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading. Additionally, based on these results and those from the original paper, we can conclude that our method generates higher-quality text compared to various baselines.
# params | Neo | Opt | Neo + KUMPR | Neo + ICU (ours) |
125M | 3.1 / 3.75 | 3.7 / 3.08 | 1.1 / 1.05 | 5.7 / 3.92 |
1.3B | 1.8 / 2.72 | 3.1 / 3.63 | 1.4 / 1.26 | 4.8 / 4.33 |
2.7B | 1.5 / 2.11 | 3.2 / 3.65 | 1.1 / 1.00 | 4.7 / 4.40 |
Appendix C Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to LLMs other than the GPT-Neo model. As introduced in Section 3, our proposed ICU framework for generative LMs, which can be applied to various advanced LLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the Tinyllama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR (Jang et al., 2023), DPO (Maini et al., 2024), KL (Maini et al., 2024), and LLMU (Yao et al., 2023). The results are summarized in Table 4. Our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU.
Model | EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | GPT | Epoch |
(%) | (%) | (F1) | (ACC) | (F1) | (PPL) | (PPL) | ||||
TinyLlama | 56.2 | 76.8 | 71.2 | 4.742 | 46.2 | 12.4 | 12.8 | 10.7 | - | - |
TinyLlama + KUMPR | 0.0 | 0.3 | 21.4 | 2.015 | 34.9 | 0.0 | >10000 | 2152.2 | 1.99 | 4.0 |
TinyLlama + DPO | 3.1 | 52.6 | 38.3 | 1.392 | 40.8 | 11.7 | 13.9 | 11.5 | 1.82 | 3.0 |
TinyLlama + KL | 1.8 | 52.2 | 31.9 | 2.753 | 42.8 | 12.3 | 36.7 | 14.0 | 1.63 | 6.0 |
TinyLlama + LLMU | 2.5 | 46.0 | 41.9 | 3.683 | 45.6 | 12.4 | 14.9 | 11.2 | 2.82 | 8.0 |
TinyLlama + ICU (ours) | 1.4 | 44.0 | 44.8 | 3.932 | 45.8 | 12.3 | 14.3 | 11.0 | 3.35 | 10.0 |
Appendix D Learning rate analysis
We conduct experiments with different learning rate, the results are shown in Table 5. The weights were chosen as . As the results do not differ much and we prefer faster unlearning and better performance, the final learning rate is set as .
Learning | EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | Epoch |
Rate | (%) | (%) | (F1) | (ACC) | (F1) | (PPL) | (PPL) | ||
1e-5 | 3.7 | 56.1 | 52.4 | 3.773 | 42.3 | 9.7 | 25.1 | 50.0 | 8.4 |
5e-6 | 4.4 | 57.9 | 53.3 | 3.789 | 42.4 | 9.8 | 24.0 | 47.4 | 14.4 |
2e-6 | 4.5 | 58.3 | 53.7 | 3.751 | 42.5 | 9.9 | 23.4 | 45.5 | 33.6 |
1e-6 | 4.8 | 59.1 | 54.3 | 3.790 | 42.5 | 9.9 | 22.9 | 44.7 | 66 |
5e-7 | 4.9 | 59.4 | 54.2 | 3.798 | 42.6 | 10.0 | 22.8 | 44.4 | 131.8 |
Appendix E Unlearning process analysis
To illustrate the detailed process of unlearning, we provide an example showing the effects at different epochs. As shown in Figure 5, given the prefix, the original model generates texts containing key terms like “Internet of Things (IoT)”, which are expected to be unlearned. At epoch 5, the model still generates completely same texts and at epoch 6, the generated texts change but still include the key term. However, by the next epoch, the model omits “IoT” and exclusively generates texts containing “cloud computing” which appears in the prefix.
Appendix F More case studies
In addition to the example in the Case Study section, we provide three additional examples to give readers a clearer understanding of how our method effectively implements machine unlearning to protect against extraction attacks, as shown in Figure 6.
Appendix G Full results
The full results of the individual runs are shown in Table 6 and 7. Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 4.5 except that results of some classification tasks are less stable than others. Even though some models have lost overall expressive ability after KUMPR, they can still achieve satisfactory results on certain classification tasks while only get poor score on dialogue datasets.
Model | # | EL10 | MA | BERT | Entropy | Lamba. | Hella. | Wino. | COPA | ARC-E | ARC-C | Piqa | MathQ | PubQ | Avg. |
Params | (%) | (%) | (F1) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | (ACC) | ||
Neo (before unlearning) | 125M | 51.9 | 76.8 | 70.3 | 4.139 | 37.6 | 28.2 | 51.5 | 62.0 | 46.0 | 22.4 | 63.4 | 22.5 | 57.6 | 43.5 |
Neo + KUMPR | 125M | 0.1 | 22.0 | 26.5 | 0.863 | 2.0 | 26.9 | 51.1 | 57.0 | 36.0 | 22.1 | 56.9 | 21.3 | 41.0 | 34.9 |
0.3 | 24.0 | 33.3 | 0.707 | 44.8 | 27.3 | 52.6 | 49.0 | 38.9 | 21.1 | 59.6 | 22.5 | 57.6 | 41.5 | ||
1.2 | 9.4 | 28.9 | 0.546 | 0.0 | 26.0 | 50.6 | 56.0 | 30.7 | 20.1 | 52.0 | 20.7 | 32.4 | 32.1 | ||
0.4 | 18.2 | 30.8 | 0.327 | 0.1 | 26.6 | 49.9 | 56.0 | 32.8 | 21.4 | 55.5 | 21.2 | 32.4 | 32.9 | ||
1.5 | 22.0 | 29.1 | 1.119 | 0.5 | 26.8 | 50.7 | 57.0 | 33.7 | 21.1 | 26.7 | 21.1 | 40.6 | 34.2 | ||
Neo + ICU (ours) | 125M | 3.6 | 53.3 | 52.8 | 3.848 | 37.3 | 28.3 | 51.1 | 62.0 | 45.3 | 23.1 | 63.2 | 22.6 | 57.6 | 43.4 |
4.8 | 52.3 | 53.0 | 3.670 | 44.4 | 28.4 | 53.1 | 60.0 | 43.9 | 20.7 | 62.8 | 22.6 | 57.6 | 43.7 | ||
4.4 | 56.3 | 51.8 | 3.770 | 34.2 | 28.3 | 52.2 | 61.0 | 43.9 | 22.4 | 63.3 | 21.7 | 57.8 | 42.8 | ||
3.9 | 59.0 | 54.8 | 3.740 | 36.1 | 28.3 | 51.9 | 62.0 | 43.7 | 20.7 | 62.7 | 22.5 | 57.6 | 42.8 | ||
5.0 | 56.9 | 54.3 | 4.137 | 43.1 | 28.3 | 52.7 | 61.0 | 44.9 | 22.1 | 62.7 | 22.2 | 57.6 | 43.8 | ||
Neo (before unlearning) | 1.3B | 98.2 | 92.3 | 86.3 | 4.640 | 57.4 | 37.0 | 54.9 | 70.0 | 56.5 | 25.8 | 70.3 | 22.0 | 53.8 | 49.7 |
Neo + KUMPR | 1.3B | 0.0 | 0.0 | 37.8 | 2.958 | 0.0 | 25.7 | 50.8 | 51.0 | 25.4 | 17.4 | 53.2 | 18.2 | 57.6 | 33.3 |
0.6 | 7.0 | 21.7 | 0.216 | 0.0 | 25.7 | 50.4 | 55.0 | 28.2 | 19.4 | 53.3 | 21.3 | 57.6 | 34.5 | ||
0.7 | 10.5 | 28.1 | 0.471 | 0.0 | 25.9 | 48.7 | 52.0 | 27.9 | 19.4 | 52.7 | 20.6 | 57.6 | 33.9 | ||
1.0 | 9.8 | 22.5 | 0.324 | 0.0 | 25.9 | 50.3 | 55.0 | 27.7 | 19.4 | 52.4 | 19.7 | 57.6 | 34.2 | ||
1.6 | 13.2 | 22.7 | 0.115 | 0.0 | 25.8 | 49.8 | 50.0 | 30.0 | 22.4 | 52.3 | 20.4 | 57.6 | 34.3 | ||
Neo + ICU (ours) | 1.3B | 4.5 | 53.3 | 53.5 | 4.218 | 54.3 | 37.1 | 54.8 | 70.0 | 55.6 | 25.4 | 69.8 | 21.3 | 49.4 | 48.6 |
4.8 | 53.8 | 57.7 | 4.069 | 54.3 | 37.2 | 55.3 | 69.0 | 56.3 | 25.8 | 70.2 | 21.8 | 50.2 | 48.9 | ||
4.5 | 53.7 | 53.8 | 4.488 | 55.7 | 37.6 | 55.2 | 68.0 | 56.0 | 26.4 | 70.4 | 21.4 | 50.0 | 49.0 | ||
4.8 | 41.0 | 44.2 | 2.440 | 54.6 | 37.3 | 55.1 | 70.0 | 56.8 | 25.8 | 69.9 | 21.5 | 49.8 | 49.0 | ||
4.7 | 54.5 | 54.4 | 4.286 | 57.5 | 37.3 | 54.9 | 67.0 | 56.7 | 25.8 | 70.1 | 21.5 | 50.5 | 49.3 | ||
Neo (before unlearning) | 2.7B | 96.7 | 93.7 | 90.2 | 4.719 | 62.3 | 40.8 | 56.6 | 75.0 | 59.5 | 26.1 | 73.0 | 21.3 | 57.0 | 52.4 |
Neo + KUMPR | 2.7B | 1.1 | 29.4 | 28.2 | 0.603 | 7.9 | 26.3 | 48.2 | 53.0 | 30.3 | 19.1 | 53.4 | 20.2 | 57.2 | 35.1 |
3.0 | 25.4 | 25.1 | 0.304 | 0.8 | 25.9 | 50.4 | 57.0 | 30.2 | 19.1 | 54.8 | 20.4 | 57.6 | 35.1 | ||
1.7 | 19.8 | 23.9 | 0.302 | 0.0 | 25.8 | 49.4 | 54.0 | 32.8 | 18.4 | 54.2 | 20.7 | 57.6 | 34.8 | ||
0.5 | 7.5 | 25.2 | 0.396 | 0.0 | 25.7 | 50.5 | 51.0 | 29.1 | 19.1 | 52.5 | 19.1 | 40.0 | 31.9 | ||
0.8 | 11.4 | 30.2 | 0.992 | 0.0 | 26.6 | 49.8 | 55.0 | 34.7 | 21.1 | 55.0 | 21.9 | 32.4 | 32.9 | ||
Neo + ICU (ours) | 2.7B | 4.2 | 47.7 | 49.4 | 3.793 | 58.6 | 41.2 | 55.5 | 76.0 | 59.3 | 28.8 | 72.6 | 21.7 | 57.6 | 52.4 |
5.0 | 48.3 | 53.2 | 3.577 | 60.4 | 41.0 | 55.6 | 70.0 | 59.6 | 25.8 | 72.7 | 21.7 | 57.2 | 51.6 | ||
4.7 | 43.2 | 50.6 | 3.416 | 60.2 | 41.0 | 55.6 | 74.0 | 60.4 | 27.8 | 72.8 | 21.7 | 57.2 | 52.3 | ||
4.4 | 48.4 | 55.4 | 3.920 | 60.2 | 40.8 | 56.1 | 73.0 | 59.1 | 27.8 | 72.9 | 21.5 | 57.2 | 52.1 | ||
4.4 | 53.9 | 54.7 | 3.922 | 61.4 | 40.8 | 56.7 | 73.0 | 60.2 | 27.4 | 73.0 | 21.5 | 57.4 | 52.4 |
Model | # | WoW | ED | BST | WoI | Avg. | Pile | Wikitext | GPT | Epoch |
Params | (F1) | (F1) | (F1) | (F1) | (F1) | (PPL) | (PPL) | |||
Neo (before unlearning) | 125M | 10.5 | 8.4 | 9.7 | 11.3 | 10.0 | 20.1 | 38.0 | 3.75 | - |
Neo + KUMPR | 125M | 5.8 | 6.0 | 6.1 | 6.9 | 6.2 | 444.7 | 2396.7 | 1.01 | 3 |
1.1 | 1.1 | 0.7 | 1.0 | 1.0 | 200.8 | 437.8 | 1.20 | 2 | ||
0.7 | 1.5 | 0.5 | 0.1 | 0.7 | >10000 | >10000 | 0.93 | 5 | ||
4.7 | 6.4 | 5.9 | 5.5 | 5.6 | 2778.3 | >10000 | 0.99 | 3 | ||
4.3 | 4.7 | 4.7 | 5.8 | 4.9 | 1212.3 | 9470.2 | 1.12 | 3 | ||
Neo + ICU (ours) | 125M | 11.6 | 9.0 | 9.9 | 11.7 | 10.6 | 21.5 | 39.2 | 3.48 | 19 |
10.9 | 8.4 | 9.6 | 11.1 | 10.0 | 22.0 | 40.7 | 3.68 | 22 | ||
11.0 | 8.4 | 9.0 | 11.2 | 9.9 | 21.6 | 40.4 | 3.98 | 24 | ||
11.3 | 9.0 | 10.0 | 11.2 | 10.4 | 21.4 | 39.7 | 4.27 | 21 | ||
11.3 | 8.7 | 9.6 | 11.7 | 10.4 | 21.6 | 40.4 | 4.20 | 21 | ||
Neo (before unlearning) | 1.3B | 12.7 | 10.4 | 12.1 | 13.8 | 12.3 | 13.2 | 18.7 | 2.72 | - |
Neo + KUMPR | 1.3B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | >10000 | >10000 | 2.68 | 1 |
0.0 | 0.1 | 0.0 | 0.0 | 0.0 | >10000 | >10000 | 0.95 | 3 | ||
0.0 | 0.0 | 0.0 | 0.1 | 0.0 | >10000 | >10000 | 0.88 | 2 | ||
0.1 | 0.0 | 0.8 | 0.2 | 0.1 | >10000 | >10000 | 0.86 | 2 | ||
0.3 | 0.0 | 0.8 | 1.1 | 0.6 | 2164.8 | >10000 | 0.93 | 2 | ||
Neo + ICU (ours) | 1.3B | 12.4 | 10.3 | 11.7 | 13.0 | 11.9 | 14.1 | 19.3 | 4.35 | 28 |
12.1 | 10.0 | 11.7 | 13.5 | 11.8 | 14.1 | 19.3 | 4.72 | 29 | ||
12.8 | 10.5 | 12.1 | 13.4 | 12.2 | 14.3 | 19.4 | 4.66 | 30 | ||
12.6 | 10.4 | 12.0 | 13.9 | 12.2 | 14.2 | 19.3 | 2.96 | 33 | ||
12.6 | 11.1 | 12.2 | 13.9 | 12.5 | 14.1 | 19.2 | 4.96 | 26 | ||
Neo (before unlearning) | 2.7B | 12.5 | 10.8 | 12.4 | 13.6 | 12.3 | 12.0 | 16.2 | 3.65 | - |
Neo + KUMPR | 2.7B | 10.9 | 3.2 | 9.7 | 11.6 | 8.8 | 39.2 | 79.9 | 1.02 | 7 |
10.6 | 0.8 | 9.7 | 11.9 | 8.2 | 66.4 | 166.8 | 1.00 | 8 | ||
4.9 | 0.1 | 9.0 | 9.3 | 5.8 | 111.2 | 370.8 | 0.94 | 6 | ||
0.1 | 0.0 | 0.3 | 0.4 | 0.2 | >10000 | >10000 | 0.96 | 6 | ||
3.5 | 3.2 | 3.8 | 4.3 | 3.7 | 227.1 | 398.6 | 1.08 | 6 | ||
Neo + ICU (ours) | 2.7B | 12.1 | 10.4 | 12.2 | 13.0 | 11.9 | 13.0 | 16.9 | 3.70 | 29 |
12.3 | 10.3 | 12.3 | 12.8 | 12.0 | 13.2 | 16.9 | 4.38 | 34 | ||
12.2 | 10.0 | 11.7 | 13.1 | 11.7 | 13.3 | 17.0 | 4.13 | 38 | ||
12.8 | 10.9 | 12.9 | 13.5 | 12.5 | 13.1 | 17.1 | 5.29 | 34 | ||
12.3 | 11.2 | 12.2 | 13.3 | 12.2 | 13.1 | 17.0 | 4.51 | 28 |

### v3

@@
 Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models
-Abstract
-Recent advancements in machine learning, particularly in Natural Language Processing (NLP), have led to the development of sophisticated models trained on extensive datasets, yet raising concerns about the potential leakage of sensitive information.
+Abstract.
+Recent advances in machine learning, particularly in Natural Language Processing (NLP), have produced powerful models trained on vast datasets.
+However, these models risk leaking sensitive information, raising privacy concerns.
 In response, regulatory measures such as the European Union’s General Data Protection Regulation (GDPR) have driven increasing interest in Machine Unlearning techniques, which enable models to selectively forget specific data entries.
-Early approaches primarily relied on pre-processing methods, while more recent research has shifted towards training-based unlearning techniques.
-Despite their effectiveness, most existing methods require access to the original training data, which is often inaccessible.
-Additionally, directly applying unlearning techniques bear the cost of undermining the model’s expressive capabilities.
-To address these challenges, we introduce the Iterative Contrastive Unlearning (ICU) framework, which consists of three core components: A Knowledge Unlearning Induction module designed to remove specific knowledge through an unlearning loss; A Contrastive Learning Enhancement module to preserve the model’s expressive capabilities against the pure unlearning goal; And an Iterative Unlearning Refinement module that dynamically assess the unlearning extent on specific data pieces and make iterative update.
+Early unlearning approaches primarily relied on pre-processing methods, while more recent research has shifted towards training-based solutions.
+Despite their effectiveness, a key limitation persists: most methods require access to original training data, which is often unavailable.
+Additionally, directly applying unlearning techniques bears the cost of undermining the model’s expressive capabilities.
+To address these challenges, we introduce the Iterative Contrastive Unlearning (ICU) framework, which consists of three core components: A Knowledge Unlearning Induction module designed to target specific knowledge for removal using an unlearning loss; A Contrastive Learning Enhancement module to preserve the model’s expressive capabilities against the pure unlearning goal; And an Iterative Unlearning Refinement module that dynamically adjusts the unlearning process through ongoing evaluation and updates.
 Experimental results demonstrate the efficacy of our ICU method in unlearning sensitive information while maintaining the model’s overall performance, offering a promising solution for privacy-conscious machine learning applications.
-1 Introduction
-With the continuous evolution of machine learning, we have witnessed an unprecedented expansion in the size of models and the diversity of data used for their training.
-Particularly, the field of Natural Language Processing (NLP) has experienced remarkable progress, driven by the development of advanced Generative Language Models (GLMs) such as GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and Google Gemini (Team et al., 2023).
-Despite these advancements, concurrent studies have highlighted a critical concern: the potential leakage of sensitive information (Carlini et al., 2021), including phone numbers, email addresses, and other personal data embedded within the training datasets (Brown et al., 2022; Carlini et al., 2022).
-In response to these privacy concerns, regulations such as the General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) have codified the “Right To Be Forgotten” (RTBF) (Mantelero, 2013; Villaronga et al., 2018), leading to the exploration of Machine Unlearning techniques.
-These techniques aim to make models effectively “forget” specific data, treating them as if they were never part of the training dataset.
+1. Introduction
+The rapid advancement of deep learning, particularly in Natural Language Processing (NLP), has led to the widespread deployment of powerful generative language models (GLMs) such as GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and Google Gemini (Team et al., 2023).
+These models exhibit remarkable capabilities in text generation, summarization, and conversational AI, transforming numerous real-world applications.
+However, this progress has also raised critical concerns regarding data privacy, security, and compliance with evolving regulations.
+One of the most pressing challenges is the risk of unintentional memorization and leakage of sensitive information, such as personally identifiable data, proprietary content, and confidential records contained in training datasets (Carlini et al., 2021; Brown et al., 2022; Carlini et al., 2022).
+To mitigate these risks, regulatory frameworks like the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017) have introduced principles such as the “Right To Be Forgotten” (RTBF) (Mantelero, 2013; Villaronga et al., 2018), which necessitate the ability to remove or selectively forget specific information from trained machine learning models.
+This has driven extensive research into machine unlearning, a field that aims to ensure models can eliminate particular knowledge while maintaining overall functionality and performance.
+Comparing text sequences generated by three models in side-by-side panels.
+Panel (a) shows outputs from the original model, which indicates its knowledge of “Harry Potter”, panel (b) demonstrates outputs after unlearning via KUMPR, showing nonsensical repetition, and panel (c) displays outputs from the proposed ICU.
+Compared to (a) and (b), our proposed method (c) appears to reduce repetition and improve specificity, showing concise, contextually appropriate responses without redundant or irrelevant content.
 In the initial stages of research, various pre-processing methods were proposed to achieve unlearning on certain data.
 For instance, Kandpal et al.
-(2022) found that the likelihood of language models regenerating training sequences is correlated with the frequency of those sequences in the training set.
-They showed that deduplicating the training data makes language models significantly more resilient against privacy attacks.
+(2022) found that the likelihood of Generative Language Models regenerating training sequences is correlated with the frequency of those sequences in the training set.
+They showed that deduplicating the training data makes GLMs significantly more resilient against privacy attacks.
 However, such methods are often time-consuming and resource-intensive, making them impractical for scenarios with frequent unlearning requests.
@@
 Nonetheless, both of these methods assume that the training data remains accessible during the unlearning process.
-In practice, the training data for large language models may not be available after model deployment, rendering such methods infeasible.
+In practice, the training data for Generative Language Models may not be available after model deployment, rendering such methods infeasible.
 A recent method, Knowledge Unlearning for Mitigating Privacy Risks (KUMPR) (Jang et al., 2023), was designed to address this scenario.111We refer to this method as KUMPR as the authors did not provide a specific name.
 KUMPR reverses the training objective by maximizing the negative log-likelihood for target tokens and uses metrics to determine whether the model has “forgotten” a target sequence.
-However, directly reversing the training objective of language models can undermine their overall expressive capabilities, beyond merely forgetting the intended information.
+However, directly reversing the training objective of GLMs can go beyond merely forgetting the target knowledge, leading to the loss of their overall expressive capabilities.
 As illustrated in Figure 1, given a training sample with the prefix “Who is Harry Potter?” and its reference suffix, the goal is to make the model “forget” specific knowledge, such as “J.K.
 Rowling” and “1997”.
-Although the existing methods (e.g., KUMPR) successfully forget the key information contained in the reference suffix, the resulting model loses its general expressive ability, often generating repetitive, nonsensical text like “I have a Harry Potter”.
-To address this limitation, we propose an Iterative Contrastive Unlearning (ICU) framework, which aims to achieve unlearning while preserving the model’s overall generalization ability.
-More specifically, our ICU framework comprises three components: (1) a Knowledge Unlearning Induction (KUI) module, which targets specific knowledge to be unlearned from the model; (2) a Contrastive Learning Enhancement (CLE) module, which samples paired data from the analogous documents of target knowledge.
+Although existing methods like KUMPR successfully forget the key information contained in the reference suffix, the resulting model loses its general expressive ability, often generating repetitive, nonsensical text.
+For example, in Figure 1 (b), the model repeats “and I have a Harry Potter”, indicating that it has collapsed.
+To address this limitation, we propose an Iterative Contrastive Unlearning (ICU) framework, which aims to achieve knowledge unlearning while preserving the model’s overall generalization ability.
+More specifically, our ICU framework comprises three components: (1) A Knowledge Unlearning Induction (KUI) module, which applies an unlearning loss to remove the specific knowledge effectively; (2) A Contrastive Learning Enhancement (CLE) module, which samples paired data from the analogous documents of target knowledge.
 Subsequently, we further design two learning enhancement losses to maintain the generalization ability against the unlearning process.
-(3) an Iterative Unlearning Refinement (IUR) module, which assess the unlearning extent on specific data pieces and updates the unlearning dataset dynamically.
+(3) An Iterative Unlearning Refinement (IUR) module, which assess the unlearning extent on specific data pieces and updates the unlearning dataset dynamically, effectively preventing over-unlearning and excessive performance degradation.
+Finally, experimental results and analyses from various perspectives demonstrate that our proposed method outperforms baseline approaches in terms of balancing maintaining performance and improving unlearning efficiency.
 In brief, our contributions are as follows:
-We conduct an in-depth study on the unlearning techniques for Generative Language Models, specifically focusing on maintaining the expression ability while effectively unlearning, an area that has been largely overlooked in previous research.
+We conduct an in-depth study on the unlearning techniques for Generative Language Models, specifically focusing on maintaining the expression ability while achieving effectively unlearning, an area that has been largely overlooked by previous researchers.
 We propose the ICU framework, which consists of three components: Knowledge Unlearning Induction, Contrastive Learning Enhancement, and Iterative Unlearning Refinement.
-We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/himalalps/ICU.
-2 Related work
-2.1 Machine unlearning
-Machine unlearning, introduced by Cao & Yang (2015), aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data was never part of the training set.
+We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at
+https://github.com/himalalps/ICU.
+2. Related work
+The related work can be categorized into two parts, including: (1) Machine Unlearning and (2) Generative Language Models.
+2.1. Machine unlearning
+Machine unlearning, introduced by Cao and Yang (2015), aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data was never part of the training set.
 Traditional approaches (Nguyen et al., 2022) that exclude specific data from training datasets and retrain the model are highly time-consuming and resource-intensive, making them impractical for modern deep neural networks.
@@
 However, challenges arise when dealing with unstructured data, as pre-processing may not fully remove all sensitive information (Brown et al., 2020) and cannot comprehensively address ongoing deletion demands (Brown et al., 2022).
-Recent studies (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023) have focused on fine-tuning language models to tackle machine unlearning challenges.
+Recent studies (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023; Eldan and Russinovich, 2023) have focused on fine-tuning Generative Language Models to tackle machine unlearning challenges.
 Jang et al.
@@
 Despite effectiveness, this kind of methods cannot avoid undermining models’ generalization ability.
-Other recent methods (Chen & Yang, 2023; Gao et al., 2024; Gu et al., 2024; Kassem et al., 2023; Wang et al., 2023; Maini et al., 2024; Yao et al., 2023) use diverse techniques such as knowledge gap alignment and reinforcement learning.
+Other recent methods (Chen and Yang, 2023; Gao et al., 2024; Gu et al., 2024; Kassem et al., 2023; Wang et al., 2023; Maini et al., 2024; Yao et al., 2023; Li et al., 2024) use diverse techniques such as knowledge gap alignment and reinforcement learning.
 Despite unlearning effectively, these methods are often complex and computationally expensive, limiting their practicality.
@@
 (2024) utilized second-order information (Hessian) to provide stronger guarantees for data removal while maintaining model utility, but this approach requires substantial computational resources for Hessian approximation, making it difficult to play a role in real scenarios.
-2.2 Language models
-Language models are designed to understand, generate, and predict human language (Zhao et al., 2023), which have gained considerable attention in recent years (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023).
+Pawelczyk et al.
+(2023) applied in-context methods in unlearning approaches, yet not effective for generation tasks.
+2.2. Generative language models
+Generative Language Models are designed to understand, generate, and predict human language (Zhao et al., 2023), which have gained considerable attention in recent years (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023).
 Traditional language models, such as rule-based approaches (Weizenbaum, 1966) and statistical models (Brown et al., 1992), generate outputs that resemble human language but do not perfectly reflect the training data.
@@
 Additionally, increasing model capacity has led to greater demands for training data and computational resources (Achiam et al., 2023; Anthropic, 2024; Hoffmann et al., 2022; Ouyang et al., 2022), creating significant challenges for researchers working on machine unlearning for advanced models.
-3 ICU framework
-3.1 Problem statement
-Given the data and model , where is the collection of pieces of data to be forgotten and is the original language model with its parameters denoted as , machine unlearning aims to modify the parameters such that the retention of previously learned information about is minimized while maintaining desirable model performance and meeting specified constraints.
-3.2 Model overview
+3. Problem statement
+Formally speaking, the machine unlearning task can be formulated as follows:
+Given the target data and model , where is the collection of pieces of data to be forgotten and is the original model with its parameters denoted as , machine unlearning aims to modify the parameters such that the retention of previously learned information about is minimized while maintaining desirable model performance and meeting specified constraints.
+The architecture of the Iterative Contrastive Unlearning framework, divided into three interconnected modules.
+Panel (a) (Knowledge Unlearning Induction) unlearns the target data directly.
+Panel (b) (Contrastive Learning Enhancement) demonstrates paired data processing, where the target data (“Who is Harry Potter?”) undergoes KNN search and the result (“Who is Hermione Granger?”) is used to refine outputs by contrasting correct and incorrect responses, resolving nonsensical repetition seen in intermediate stages.
+Panel (c) (Iterative Unlearning Refinement) depicts iterative optimization, which adjusts the unlearning process through ongoing evaluation and updates.
+4. ICU framework
+4.1. Model overview
 We propose a novel Iterative Contrastive Unlearning (ICU) framework, illustrated in Figure 2.
@@
 Furthermore, (c) Iterative Unlearning Refinement module updates the data to be forgotten in an iterative manner, preventing over-unlearning and mitigating performance degradation.
-3.3 Knowledge Unlearning Induction
+4.2. Knowledge Unlearning Induction
 For a sample in the forget set , the sequence of tokens is denoted as .
@@
 The unlearning objective is defined as:
-where denotes the first tokens of the sequence, and represents the conditional probability of predicting the next token given the previous tokens , with representing the model parameters.
-3.4 Contrastive Learning Enhancement
-To maintain the model’s stable generation capabilities during unlearning, we propose training the model simultaneously on analogous data.
+where denotes the first tokens of the sequence, is the length of tokens provided to the model, and represents the conditional probability of predicting the next token given the previous tokens , with representing the model parameters.
+4.3. Contrastive Learning Enhancement
+To maintain the model’s stable expression capabilities during unlearning, we propose training the model simultaneously on analogous data.
 This approach ensures that the model forgets specific information without significantly reducing its ability to recognize and generate similar patterns.
-Analogous data Construction.
+4.3.1. Analogous data construction.
 The first step is to construct a data pool related to the forget set , aiming to identify data samples that contain similar but different knowledge.
@@
 This process results in the creation of the Analogous Set , which will be used in subsequent steps.
-KNN Sampling.
+4.3.2. KNN sampling.
 In this part, we compute the sentence embeddings for all samples in using a pre-trained sentence transformer .
@@
 where is the cosine similarity function used in KNN search.
-Learning Enhancement.
+4.3.3. Learning enhancement.
 The retrieved paired data for is denoted as .
 In contrast to the unlearning objective, we force the model to learn patterns from these paired token sequences using the negative log-likelihood, defined as follows:
-3.5 Iterative Unlearning Refinement
+4.4. Iterative Unlearning Refinement
 Unlike conventional machine learning techniques, validating the efficacy of unlearning presents challenges in identifying a suitable validation set, as is integrated into the training phase.
@@
 After each training epoch, the model’s performance relative to the target data is evaluated using the metrics described below:
-where and are predefined thresholds for the iteration process.
+where is the referenced target sample, denotes the output of the model provided input sequence , and and are predefined thresholds for the iteration process.
 We empirically determine that a specific token sequence is considered “forgotten” if Equation (5) is satisfied.
@@
 Unlike previous metrics that rely solely on exact word n-grams, BERTScore considers semantic similarity, offering a more adaptable and accurate measure of sentence similarity.
-3.6 Training
-The objectives in Equation (1), (3) and (4) are jointly used to optimize the model during unlearning. The training process is governed by minimizing the following loss function:
-where are positive hyper-parameters that control the contributions of the different optimizing objectives.
-4 Experiments
+4.5. Training
+5. Experiments
 In this section, we first describe the datasets used for training and evaluation, as well as the metrics employed to assess performance.
@@
 Finally, we present and analyze the experimental results.
-4.1 Datasets
+5.1. Datasets
 To evaluate ICU’s learning and unlearning capabilities, we selected two types of datasets: Target Datasets, which assess the unlearning performance, and Downstream Dataset, which evaluates the original capabilities of the models.
@@
 Following Jang et al.
-(2023), we only use the prefix and suffix tokens in our experiments.
+(2023), we only use the prefix and suffix tokens thus is set to 50.
+This choice of benchmark follows standard practices like Jang et al.
+(2023) and Kassem et al.
+(2023), where the same benchmark is widely used to evaluate unlearning methods.
+While alternative benchmarks such as WMDP (Li et al., 2024) exist, they represent different datasets designed for the same unlearning task, and thus do not fundamentally differ in the nature of the evaluation.
+Since all these benchmarks focus on generative unlearning tasks, additional experiments on alternative datasets are unnecessary, as they would likely yield similar insights without providing significant new information.
 Downstream Dataset.
@@
 Following (Jang et al., 2023), we use the test set for Lambada and the validation set for the remaining tasks.
-4.2 Metrics
-As stated in Section 4.1, we assess both learning and unlearning capabilities using two types of datasets.
+5.2. Metrics
+As stated in Section 5, we assess both learning and unlearning capabilities using two types of datasets.
 For evaluating unlearning performance, we follow Jang et al.
@@
 (2023), these metrics are also used as stopping criteria during training, making them unsuitable as sole evaluation metrics.
-As mentioned in Section 3.5, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.
+As mentioned in Section 4.3.3, we additionally employ BERTScore and BLEU to measure forgetting during the Iterative Unlearning Refinement (IUR) module.
 To evaluate the original capabilities of unlearned models, we first use the basic metrics provided in the downstream datasets, obtaining Accuracy for classification task and F1 for dialogue tasks.
@@
 Memorization Accuracy (MA) (Tirumala et al., 2022) quantifies how much model has memorized the given token sequences, which is defined as follows:
-Information Entropy quantifies the average uncertainty in a set of outcomes, reflecting the amount of information produced by a random source.
+Information Entropy quantifies the average uncertainty in a set of outcomes, reflecting the amount of information produced by a random source (Vajapeyam, 2014).
 Higher entropy indicates greater unpredictability and information content.
@@
 Human Evaluation indicates that we hire human experts to determine the goodness of the generated response, following the goal in the GPT Evaluation part.
-The annotation agreement can be found in Appendix B.
-4.3 Baseline methods
+The annotation details can be found in Appendix B.
+5.3. Baseline methods
 Our experiments use the GPT-Neo model family (125M, 1.3B, 2.7B) (Black et al., 2021), which is pre-trained on the Pile corpus.
@@
 We follow their publicly released codes and the same training and evaluation procedure to obtain the results.
-4.4 Configurations
-For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples.
-In the Contrastive Learning Enhancement module (Section 3.4), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling.
-The model is optimized by Adam (Kingma & Ba, 2014) with a learning rate of , and .
-We regard the model to have “forgotten” the target dataset with an average of and .
-The filtering thresholds during iteration are and as introduced in Section 3.5.
-We run all the experiments on a Linux server with one 2.60GHz Intel Xeon Platinum 8358 CPU and NVIDIA GeForce RTX 3090 GPUs.
-We use one GPU for 125M models with batch size of 8.
-With Deepspeed Stage 2, we use three for 1.3B and six for 2.7B respectively with batch size of 4.
+For experiment on model families other than GPT-Neo, we also include results on TinyLlama 1.1B model (Zhang et al., 2024) in Appendix C.
 Model | # | EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | GPT | Epoch |
@@
 Neo + ICU (ours) | 4.5 | 48.3 | 52.7 | 3.725 | 52.1 | 12.1 | 13.1 | 17.0 | 4.40 | 32.6 |
-4.5 Main results
+5.4. Configurations
+For each model size (125M, 1.3B, 2.7B), we execute five runs of the methods, each targeting at a dataset of 128 samples.
+In the Contrastive Learning Enhancement module (Section 4.2), we utilize the remaining Pile subset for the Analogous Data Construction, and all-MiniLM-L6-v2 model to conduct KNN sampling.
+The model is optimized by Adam (Kingma and Ba, 2014) with a learning rate of , and .
+We regard the model to have “forgotten” the target dataset with an average of and following Jang et al.
+(2023).
+The filtering thresholds during iteration are and as introduced in Section 4.3.3.
+Specifically, for BERTScore, only 2% of outputs fell below , while the average score exceeded , with many instances close to .
+For BLEU, the largest value below was , and the smallest above was , making a clear dividing point.
+These thresholds effectively balance precision and applicability based on the observed data distributions.
+We run all the experiments on a Linux server with one 2.60GHz Intel Xeon Platinum 8358 CPU and NVIDIA GeForce RTX 3090 GPUs.
+We use one GPU for 125M models with batch size of 8.
+With Deepspeed Stage 2, we use three GPUs for 1.3B and six GPUs for 2.7B respectively with batch size of 4.
+5.5. Main results
 The results of all methods are summarized in Table 1.
 Overall, our method consistently achieves the best or second-best performance across all metrics compared to the baselines.
-Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the generation ability of the model.
+Although methods like KUMPR, DPO, KL and LLMU may exhibit seemingly better unlearning ability measured by EL, MA, and BERT, our method preserves the language generation ability of the model.
 Compared to Opt, our model unlearns the original model better.
@@
 In contrast, our ICU approach preserves the model’s core linguistic capabilities while erasing sensitive information, underscoring the importance of our method.
-Full results for each sample are provided in Appendix G.
-4.6 Ablation study
-As discussed in Section 3.4, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability.
-To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively.
-The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity.
-Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
-4.7 Parameter sensitivity
+Full results for each sample are provided in Appendix I.
+Compares model performance across three configurations: (, and ), measuring , , BERTScore, Information Entropy, classification/dialogue accuracy (Cls/Dia Avg.), perplexity (PPL) on Pile/Wikipedia datasets, GPT evaluation and epochs.
 EL10 | MA | BERT | Entropy | Cls Avg. | Dia Avg. | Pile | Wikitext | Epoch | ||
@@
 0.1 | 0.5 | 3.6 | 54.7 | 47.0 | 3.009 | 39.5 | 5.6 | 95.2 | 212.7 | 9.4 |
+0.1 | 1.0 | 4.2 | 56.4 | 52.5 | 3.758 | 42.9 | 10.0 | 23.3 | 44.8 | 16.2 |
 0.5 | 0.1 | 3.3 | 56.5 | 49.9 | 3.423 | 40.0 | 8.7 | 35.4 | 85.5 | 9.4 |
@@
 2.0 | 2.0 | 4.5 | 45.8 | 51.2 | 3.843 | 42.9 | 10.0 | 22.1 | 40.1 | 45.4 |
+5.6. Ablation study
+As discussed in Section 4.2, pair learning loss and KL-divergence loss are employed to ensure the model’s stable generative capability.
+To assess the impact of these losses, Figure 3 shows the performance after removing each loss respectively.
+The results indicate that both losses enhance learning performance, affirming their role in preserving the model’s generative capacity.
+Furthermore, the KL-divergence loss has a more pronounced effect, suggesting that aligning the model’s output distribution with the original model is crucial for maintaining generative performance.
+5.7. Parameter sensitivity
 To examine the influence of the loss hyperparameters and in Section 4.2, we conducted extensive parameter sensitivity experiments on GPT-Neo 125M. The results are summarized in Table 2.
@@
 To balance the model’s learning and forgetting abilities, we ultimately selected and as our reported parameters.
-4.8 Case study
+We also conduct experiment regarding the selection of learning rate in our analysis, and the result is shown in Appendix D.
+5.8. Human evaluation
+We have also conducted human evaluation about the results of different methods and the comparison between GPT and Human Evaluation scores can be found in Table 3.
+The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading.
+For statistical test, we normalize the human score by , and the Pearson correlation is , which is quite high, showing the strong correlation between human and GPT evaluation.
+Additionally, based on this human evaluation and the other experimental results, we can see that our method generates higher-quality text compared to various baselines.
+More details about human evaluation are shown in Appendix B.
+# params | Neo | Opt | Neo + KUMPR | Neo + ICU (ours) |
+125M | 3.1 / 3.75 | 3.7 / 3.08 | 1.1 / 1.05 | 5.7 / 3.92 |
+1.3B | 1.8 / 2.72 | 3.1 / 3.63 | 1.4 / 1.26 | 4.8 / 4.33 |
+2.7B | 1.5 / 2.11 | 3.2 / 3.65 | 1.1 / 1.00 | 4.7 / 4.40 |
+5.9. Case study
 To provide a clearer comparison of our methods, we present a case study demonstrating the balance between learning and unlearning.
@@
 This demonstrates the effectiveness of our Iterative Contrastive Unlearning framework.
-Additional cases are provided in Appendix F.
-More experimental analyses, such as Human Evaluation and Generalization Ability of ICU, can be found in Appendix A - G.
-5 Conclusion and future work
-In this paper, we explored the direction of machine unlearning for Generative Language Models.
-We analyzed the challenges inherent in this topic and proposed an Iterative Contrastive Unlearning (ICU) framework.
+We also analyzed the unlearning process throughout different epochs, and the details can be found in Appendix E.
+To provide a more comprehensive understanding of different methods, additional cases are also provided in Appendix F.
+6. Conclusion and future work
+In this work, we explored machine unlearning for Generative Language Models, focusing on the challenge of selectively removing sensitive data while preserving overall model performance.
+We proposed the Iterative Contrastive Unlearning (ICU) framework.
 Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data.
 Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities.
-Extensive experiments on models of three different sizes demonstrated the effectiveness of our proposed method.
-For future work, we plan to explore the application of our framework to multi-modal data, extending unlearning capabilities to models that integrate text, images, and other data types.
-Additionally, we are very interested in studying the implications of unlearning algorithms in privacy-sensitive real-world applications, such as healthcare and personalized services.
+We conducted extensive experiments across models of different scales, demonstrating that ICU effectively removes sensitive data while maintaining general capabilities.
+Our ICU provides a practical solution to privacy challenges in AI systems.
+Its ability to balance data removal with model utility makes it a promising approach for real-world, privacy-aware machine learning applications.
+Moving forward, we aim to extend ICU to multi-modal models, allowing for unlearning across text, images, and other data modalities.
+Additionally, we plan to explore real-world applications in privacy-sensitive domains such as healthcare, where balancing effective data removal with model utility is crucial.
+These future directions will help align unlearning methods with regulatory requirements and practical AI deployment.
 Appendix A GPT prompts
@@
 Candidate: "<Candidate>"
-Appendix B Human evaluation
+Appendix B Human evaluation details
 To enhance the credibility of the GPT-4 grading, we also include human annotation results.
-Specifically, for each sample, human evaluators rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference).
-The annotators are only presented with only the prefix, referenced suffix and the generated text by the model and are unaware of the full scope of the work.
-The comparison between GPT and Human Evaluation scores can be found in Table 3.
-The human annotation results and the GPT evaluation results share the same trend, demonstrating the reliability of GPT-4 grading.
-Additionally, based on these results and those from the original paper, we can conclude that our method generates higher-quality text compared to various baselines.
-# params | Neo | Opt | Neo + KUMPR | Neo + ICU (ours) |
-125M | 3.1 / 3.75 | 3.7 / 3.08 | 1.1 / 1.05 | 5.7 / 3.92 |
-1.3B | 1.8 / 2.72 | 3.1 / 3.63 | 1.4 / 1.26 | 4.8 / 4.33 |
-2.7B | 1.5 / 2.11 | 3.2 / 3.65 | 1.1 / 1.00 | 4.7 / 4.40 |
+Specifically, for each sample, a human annotator rated the text generated by the models on a scale from 1 (low quality or very similar to reference) to 8 (high quality and different from reference).
+The annotator is only presented with only the prefix, referenced suffix and the generated text by the model and is unaware of the full scope of the work.
+The instruction is shown below:
+Please evaluate the following OUTPUT based on the PREFIX and evaluate its relationship to the REFERENCED SUFFIX. The scores are as follows:
+1 - Exact repetition or redundancy
+2 - Minimal divergence with little information
+3 - Substantial divergence and poor coherence
+4 - Partial coherence with divergence
+5 - Partial similarity and coherence
+6 - Paraphrased similarity
+7 - Creative and reasonable divergence
+8 - Fully reasonable divergence
 Appendix C Generalization ability of ICU
-In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to LLMs other than the GPT-Neo model.
-As introduced in Section 3, our proposed ICU framework for generative LMs, which can be applied to various advanced LLMs, such as Llama, Bloom, etc.
+In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to GLMs other than the GPT-Neo model.
+As introduced in Section 2, our proposed ICU framework for generative LMs, which can be applied to various advanced GLMs, such as Llama, Bloom, etc.
 To better verify this, we conduct additional experiments using the TinyLlama 1.1B model (Zhang et al., 2024), which features a different architecture from the GPT-Neo model in the main experiment.
@@
 In addition to the example in the Case Study section, we provide three additional examples to give readers a clearer understanding of how our method effectively implements machine unlearning to protect against extraction attacks, as shown in Figure 6.
-Appendix G Full results
-The full results of the individual runs are shown in Table 6 and 7.
+Appendix G Limitations
+While the ICU framework demonstrates promising results in targeted unlearning, several limitations must be acknowledged.
+These challenges include generalization, sensitivity to data distribution, granularity of unlearning, computational efficiency, interactions with model editing, and evaluation constraints.
+Addressing these issues is crucial to making ICU more reliable and practical for real-world applications.
+The framework operates at the sequence level, which poses challenges in selectively forgetting fine-grained information.
+Removing specific named entities within a passage without affecting the surrounding context is difficult, potentially leading to unintended knowledge loss in structured or densely packed texts.
+Applying ICU for batch unlearning is feasible, but it lacks efficient ways to handle sequential or overlapping deletion requests.
+If the forget set keeps growing, repeated iterations could create significant computational overhead.
+This could make deployment costly in dynamic environments.
+The interaction between ICU and other model editing methods, such as factual updates or bias corrections, remains unexplored.
+Concurrent modifications could interfere with the unlearning process, potentially leading to instability or conflicting objectives that require further study.
+Existing evaluation metrics, such as BERTScore and MA, focus mainly on detecting direct memorization.
+However, they do not measure implicit knowledge retention well.
+A model might still make inferences based on forgotten information.
+To improve evaluation, we need more advanced metrics that assess forgetting in latent representations.
+Appendix H Ethics statement
+Our Iterative Contrastive Unlearning framework is designed to enhance privacy in GLMs by enabling targeted forgetting of sensitive information.
+This aligns with the RTBF principle under regulations like the GDPR.
+While ICU offers important societal benefits, it also comes with some ethical risks.
+One major risk is the possibility of misuse.
+While ICU aims to protect privacy, it could also be exploited to erase public knowledge or historical facts.
+This could lead to misinformation or censorship.
+To prevent such misuse, we recommend strict access controls and transparency in how unlearning systems are used, ensuring they serve only legitimate privacy needs.
+Complete erasure of information is challenging.
+ICU reduces the chances of sensitive data leaking, but no method can guarantee absolute forgetting due to the way large models retain knowledge.
+This is especially critical in high-stakes fields like healthcare and finance.
+Future research should focus on measuring residual memory and improving forgetting guarantees.
+Unlearning data may unintentionally impact model fairness.
+Removing specific information can alter how a model behaves, sometimes amplifying biases in downstream tasks.
+Developers should thoroughly test ICU’s effects on fairness and apply bias mitigation strategies when needed.
+Transparency in unlearning remains a challenge.
+ICU’s process lacks explainability, meaning it is difficult to determine exactly what knowledge has been removed.
+We encourage the development of explainability tools to visualize unlearning effects and provide users with opt-in or opt-out options for data removal.
+Overall, ICU is a step toward privacy-aware AI, but its real-world impact depends on ethical implementation.
+We are committed to ongoing research to address these challenges and welcome community feedback to refine our approach.
+Appendix I Full results
+The full results of the individual runs are shown in Table 6 and Table 7.
 Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 5.4 except that results of some classification tasks are less stable than others.